OpenStudy (anonymous):

Twenty-two is what percent of 55? A. 50% B. 40% C. 22% D. 250%

OpenStudy (anonymous):

22/55=0.4*100=40% OR B

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Stop helping them right off the bat. Next time let them try themselves first. We have shown them many times how to do this kind of problem.

OpenStudy (anonymous):

true true...